OpenStudy (anonymous):

Solve: negative 1 over 2 x + 10 = −x + 12

OpenStudy (caozeyuan):

so you mean\[\frac{ -1 }{ 2x}+10=-x+12\]

OpenStudy (caozeyuan):

multiply both sides by x\[\frac{ -1 }{ 2 }+10x=-x ^{2}+12x\]

OpenStudy (caozeyuan):

I think you can solve it now

OpenStudy (texaschic101):

or is it this : (-1/2)x + 10 = -x + 12 ???